# Cosmology with voids

### v3

@@ 4.1 Maximum likelihood and MCMC
$\chi^{2}=\sum_{i=1}^{n_{v}}\left[y_{i}^{\text{(measured)}}-y_{i}^{\text{(theory)}}\right]^{2}/\sigma_{i}^{2}$ (35)
@@
to keep numerical values close to unity. The theoretical values, $y_{i}^{\text{(theory)}}$  come from using the measured depths $\varepsilon_{i}$ in equation (30) with a given set of parameters $\Omega_{\text{M}}$, $\Omega_{\Lambda}$ and $h$. The “uncertainties” $\sigma_{i}$ in the expression for $\chi^{2}$ are
$\sigma_{i}^{2}=\sigma_{y,i}^{2}+\frac{f^{2}\eta^{2}}{9}\sigma_{\varepsilon,i}^{2},$ (37)
where the $\sigma_{y}$ and $\sigma_{\varepsilon}$ correspond to noise levels in outflow and depth respectively. The quantities $f$ and $\eta={H_{\text{eff}}}/{H_{\text{ref}}}$ are determined for a given model at the redshift of the $i^{\text{th}}$ void. The noise model adopted here is that measurements have normally distributed fractional errors at a level $e_{\text{rel}}\sim 0.001$–0.1. By including the noise at some tunable level, we can identify correlations in fitted parameters that are clearly the result of the cosmology and not numerical artifacts or our use of an approximation for theoretical predictions. While not our focus here, the addition of tunable noise also allows us to explore the impact of observational errors.
We use a maximum-likelihood approach to fit measurements with model parameters. In this framework, a likelihood function $\cal{P}$ returns the probability of obtaining specific measurements ($\varepsilon$, ${H_{\text{eff}}}$) given a set of model parameters $\Omega_{\text{M}}$, $\Omega_{\Lambda}$, and $h$ (the curvature contribution is derived from $1=\Omega_{\text{M}}+\Omega_{\Lambda}+\Omega_{\text{K}}$). Fitting measurements with models amounts to sampling and maximizing $\cal{P}$ in parameter space. We use the emcee Markov-chain Monte Carlo (MCMC) package (Foreman-Mackey et al., 2013) to do the sampling, with multiple “walkers” mapping out the probability distribution $\cal{P}$ in the domain $\Omega_{\text{M}}=[0.045,1]$, $\Omega_{\Lambda}=[0,1]$, $h=[0.5,0.9]$. We use uninformative priors (all models in this domain are equally likely). We thus distribute the starting location of the walkers in parameter space with a low-discrepancy quasirandom number generator (scipy.qmc). We determine best-fits from the global maximum of the likelihood function as mapped by the MCMC walkers; we obtain uncertainties and correlations between parameters from the distribution of samples in parameter space.
@@
Figure 8: A corner of corner plots (Foreman-Mackey, 2016) showing probability distributions for parameter fits to void depth and outflow rates. The labels indicate the number of voids measured. The lines locate the true values of the parameters $h$ (orange), $\Omega_{\text{M}}$ (turquoise), and $\Omega_{\Lambda}$ (blue). The fits are most sensitive to $h$, then $\Omega_{\text{M}}$, and finally $\Omega_{\Lambda}$. In the upper corner plot, showing likelihood distributions for a single void, only the Hubble parameter is well constrained. Asserting prior knowledge of $h$ at its true value selects the probability distribution in a slice along the orange lines. With that prior, the (posterior) probability distribution for $\Omega_{\text{M}}$ becomes sharply peaked around its true value. In the lower left corner plot (two voids, $z=0$ and $z=1$), both $h$ and $\Omega_{\text{M}}$ are well constrained. In the lower right panel, with three voids ($z=0,0.5,1$), $\Omega_{\Lambda}$ is also well constrained.
@@ 4.2 The redshift range
The Hubble diagram built from Type Ia supernovae provides better constraints on cosmological parameters when applied over a large redshift range (Riess et al., 1998; Perlmutter et al., 1999). In analogy voids provide better constraints when we measure $fH$ as a function of $z$. Figure 9 illustrates this effect based on samples of 100 voids spaced uniformly in $z$. The broadest range of redshifts ($z=0$–2) yields the tightest constraints..
@@
There are two main challenges for measuring voids properties over a wide redshift range. At low redshift, the observable volume available to sample voids is small; in an absurd limit, there can be at most only one void of radius 100 $h^{-1}$pc within a survey of that depth. At high redshift the challenge is measurement of void properties from faint and distant galaxies. For this theoretical demonstration of the connection between void parameters and the underlying cosmology, we consider only the first challenge, identifying the number of voids we expect as a function of redshift.
Figure 10 shows the number of voids with comoving radii around 30 $h^{-1}$pcin redshift bins from $z=0$ to $z=2$ with $\Delta z=0.1$ in a $\Lambda$CDM universe. At the lowest redshift, with $z\leq 0.1$, the survey volume is a sphere of radius of about 300 $h^{-1}$pc. As in the figure, we expect 50 voids in that region. This number count matches our expectation for 30 $h^{-1}$pc voids that represent the least dense (greatest depth) quartile of independent samples of the matter density field smoothed at that scale. We assume that this tail of the void depth distribution represents detectable voids, even at $z>0$, and that the comoving space density of voids remains constant for the redshifts we consider. As we move from $z=0$ to higher redshift, the volume sampled in a redshift bin grows significantly. At $z=0.3$ and above we anticipate a thousand voids or more per bin. The dependence of comoving distance on redshift limits the number of anticipated voids to less than $10^{4}$ per bin at higher redshift. At a redshift of $\sim$2.5, the proper volume per bin, as well as the number of voids, begins to fall with increasing redshift. If we set a depth threshold in our definition of a void that is independent of redshift, then we lose some voids at redshift $z>0$ because they are less evolved. We estimate that with a threshold of $\sim$0.2, we still retain over a thousand $\sim$30 $h^{-1}$pc per bin for $0.4\lesssim z\lesssim 2$. The bins with the most voids (over 2,000) are at redshifts of $z\approx 1$.
@@ 4.3 Summary of cosmology from idealized voids
By measuring void depth $\varepsilon$ and outflow rate ${H_{\text{eff}}}$ in voids at multiple redshifts, we build samples of $fH$ that provide constraints on $h$, $\Omega_{\text{M}}$ and $\Omega_{\Lambda}$. This procedure mimics the approach to cosmological constraints based on Type Ia supernovae (Kirshner, 1999). Preliminary tests with fits to measurements from solutions to full nonlinear shell model in section 3.3 indeed suggest that samples of void properties over a range of redshifts provide robust cosmological parameter estimates.
@@ 5 Voids in numerical realizations of the cosmic web
We create realizations of large chunks of the universe with an $n$-body particle algorithmm to evaluate the cosmological information encoded by voids. Our approach is based on the adhesion approximation (Weinberg & Gunn, 1990); similar $n$-body realization methods include the PATCHY code (Kitaura et al., 2014). In our implementation, we use the Zel’dovich approximation in step-wise fashion, allowing particles to free stream with updates as large-scale structures emerge. Critically, the adhesion approximation adds small-scale viscous diffusion to mitigate the problem of trajectory crossings that arise when free streams of matter in the Zel’dovich flows pass right through one another. As in reference (Weinberg & Gunn, 1990), this feature limits the spatial resolution of the realizations to about a megaparsec. Appendix A summarizes the implementation.
We adopt the adhesion approximation primarily because it is a fast, accurate way to track void dynamics. Because density fluctuations within voids are low amplitude, nonlinear dynamical effects are weak. Recent work (Hamaus et al., 2015; Stopyra et al., 2021; Schuster et al., 2023) demonstrates that voids are well suited to the Zel’dovich approximation and the related adhesion algorithm. The realizations of large-scale structure we construct are geared toward toward modelling voids in the linear and translinear regimes; they are distinct from full $n$-body simulations that track small-scale nonlinear dynamics with Poisson solvers and ODE integrators.
@@ 5.1 Finding voids
A key step in assessing the connection between voids and their background cosmology is finding the voids. This process is complex mainly because the cosmic web is not Swiss cheese or even spongelike, with nicely spherical holes embedded in a high-density background. Instead, choices about density thresholds (voids are not vacuums), sizes and shapes are necessary and lead to a plethora of distinct void-finding algorithms, (e.g., Kauffmann & Fairall, 1991; El-Ad & Piran, 1997; Sheth & van de Weygaert, 2004; Platen et al., 2007; Neyrinck, 2008; Sutter et al., 2015; Cai et al., 2015; Nadathur & Hotchkiss, 2015a, b; Shim et al., 2023; Zaidouni et al., 2024).
@@
We choose to implement our own void finding algorithm so that we could most easily assess the impact of changes in the algorithm parameters, balancing code speed, simplicity, and effectiveness at identifying comparatively shallow voids that fit with our analysis. Other options include the adoption of publicly available void finders (Villaescusa-Navarro, 2018; Paz et al., 2023).
@@ 5.2 Cosmological parameter estimation
The theoretical curves in figure 14 are fingerprints of the background cosmology. To extract cosmological parameters by fitting the $n$-body samples shown in figure 14 with predictions from various cosmological models, we use the same MCMC-based parameter estimators as in section 4: We seek values in a parameter space of $\Omega_{\text{M}}$, $\Omega_{\Lambda}$ and $h$ using a $\chi$-square measure to assess the goodness of fit between theory (equation (30)) and samples of $(\varepsilon,{H_{\text{eff}}})$ from voids in the $n$-body realizations. Our priors are uniform inside the domain $1.050<\Omega_{\text{M}}<$ 1.05, and $-0.1<\Omega_{\Lambda}<1$, expanded slightly from the analysis of idealized voids to better explore the sensitivity of the estimators. The scatter in the samples used arises from deviations of the $n$-body realization voids from the ideal ones.
@@
Figure 15: Corner of corners plot (Foreman-Mackey, 2016) showing MCMC samples in parameter space for independent realizations of the $\Lambda$CDM model at redshifts from $z=0$ to $z=2$. Each realization lies in a cube of 500 $h^{-1}$pc to a side and has $n=512^{3}$ points. Voids of radii $30\pm 5$ $h^{-1}$pc are included. As in figure 9 for the ideal voids, the panels illustrate that a broader range of redshifts among void samples enhances parameter estimation.
@@ 5.3 The largest void in the universe
The distribution of void sizes can provide a powerful test of the background cosmology (Pan et al., 2012). Here we focus on “the largest void”. The existence of extremely large voids may place strong constraints on the background model parameters (Blumenthal et al., 1992). The void identification criteria for this investigation are only slightly more stringent than finding local minima in tophat-smoothed density fields (Szapudi et al., 2015). Voids in our catalog have profiles that are deep in their central region, and rise to a half-depth value at the nominal radius, behavior that requires specific phase alignment between density fluctuation modes. Supervoids — composed of multiple smaller voids — might be missed by our algorithm, depending on their substructure.
@@
Figure 16 shows an example of large voids within lightcone realization of 5000 $h^{-1}$pc to a side ($n=1024^{3}$). Because of grid effects, density modes are poorly sampled for voids with radii $\gtrsim$250 $h^{-1}$pc. All depths are scaled to the present epoch. For comparison, we also show the Eridanus Supervoid parameters (Szapudi et al., 2015). The 1-$\sigma$ errors for the supervoid suggest that it is consistent with voids predicted in the $\Lambda$CDM realization. Since its initial discovery, the supervoid region has been reanalyzed (Mackenzie et al., 2017; Kovács et al., 2022). The supervoid is estimated to have a greater depth and smaller radial extent, but it is still consistent with standard theory.
Figure 16: Void depth versus size of the largest voids in a 5000 $h^{-1}$pc lightcone realization. This depth-size diagram applies to voids larger than 100 $h^{-1}$pc, scaled according to linear theory. The depth is effectively at the present epoch. The darker solid curve is $\sigma_{\text{R}}$ from the linear power spectrum; we eliminate void candidates with amplitudes below this value. The lighter-tone line is the amplitude of a one-of-a-kind void predicted in linear theory for the survey volume. The color of each sample point indicates whether the void is a “subvoid” (orange) or a void that is not embedded in a larger void (blue). The error bars show the observed Eridanus supervoid (Szapudi et al., 2015).
@@ 6 Voids in redshift space
So far, we have used the positions and peculiar velocities of galaxies as tracers to constrain the background cosmological parameters (section 5.2). However, observations do not yet measure accurate distances. Instead, extensive galaxy surveys provide the distribution of objects in redshift space. Here we consider the cosmological information content of voids as mapped in redshift space (Paz et al., 2013; Pisani et al., 2015; Cai et al., 2016; Nadathur et al., 2019a, 2020b; Correa et al., 2021, 2022; Hamaus et al., 2022; Woodfinden et al., 2022, 2023).
@@
$\vec{s}$$\approx$ $\vec{S}+H\vec{r}+\hat{e}_{\text{los}}\cdot\vec{v}\hat{e}_{\text{los}}$ (38)
$\approx$ $\vec{S}+H\left(\vec{r}+\frac{f\varepsilon\Delta_{\text{los}}}{3}\hat{e}_{\text{los}}\right)$ (39)
@@
The mapping between real space and redshift space significantly impacts the measurement of void properties (Kaiser, 1987). A small region in real space within an idealized void is stretched in redshift space along the line-of-sight by a factor of ${H_{\text{eff}}}/H$. The void thus appears deeper in redshift space:
$\varepsilon_{s}=\left(1+\frac{f}{3}\right)\varepsilon$ (40)
@@
Void shapes are also affected by redshift space distortions (Kaiser, 1987). However, the net fractional elongation along the line-of-sight tends to be small because the outflow velocities are small compared with the Hubble flow across the void diameter. In the ideal case, we define the shape of a void in redshift space using the location of particles at the interface with the background universe. Neglecting any shell crossings, we approximate the void shape in redshift space as a prolate ellipse, using the ellipticity,
$e_{f}={H_{\text{eff}}}/H-1\approx\frac{f\varepsilon}{3},$ (41)
@@
Successful strategies to tease out the small-amplitude distortion of void geometry in redshift space focus on combining data from multiple voids. Aggregate measures like the void-galaxy correlation function quantify averaged redshift-space distortions that emerge away from the void centers, revealing peculiar motion (Paz et al., 2013). For idealized voids, this information corresponds to the mean ellipticity at the void boundary. When calibrated with numerical simulations, the void-galaxy correlation function is impressively effective as a cosmological indicator (Nadathur et al., 2020b; Hamaus et al., 2022). Outcomes include a measure of dark energy with the Alcock-Paczynski test and the velocity growth parameter $f$ (e.g., Nadathur et al., 2020b; Hawken et al., 2020; Woodfinden et al., 2022; Wilson & Bean, 2023).

### v4

@@ 4.1 Maximum likelihood and MCMC
$\chi^{2}=\sum_{i=1}^{n_{v}}\left[y_{i}^{\text{(measured)}}-y_{i}^{\text{(theory)}}\right]^{2}/\sigma_{i}^{2}$ (4.1)
@@
to keep numerical values close to unity. The theoretical values, $y_{i}^{\text{(theory)}}$  come from using the measured depths $\varepsilon_{i}$ in equation (3.16) with a given set of parameters $\Omega_{\text{M}}$, $\Omega_{\Lambda}$ and $h$. The “uncertainties” $\sigma_{i}$ in the expression for $\chi^{2}$ are
$\sigma_{i}^{2}=\sigma_{y,i}^{2}+\frac{f^{2}\eta^{2}}{9}\sigma_{\varepsilon,i}^{2},$ (4.3)
where the $\sigma_{y}$ and $\sigma_{\varepsilon}$ correspond to noise levels in outflow and depth respectively. The quantities $f$ and $\eta={H_{\text{eff}}}/{H_{\text{ref}}}$ are determined for a given model at the redshift of the $i^{\text{th}}$ void. The noise model adopted here is that measurements have normally distributed fractional errors at a level $e_{\text{rel}}\sim 0.001$–0.1. By including the noise at some tunable level, we can identify correlations in fitted parameters that are clearly the result of the cosmology and not numerical artifacts or the result of our use of an approximation for theoretical predictions. While not our focus here, the addition of tunable noise also allows us to explore the impact of observational errors.
We use a maximum-likelihood approach to fit measurements with model parameters. In this framework, a likelihood function $\cal{P}$ returns the probability of obtaining specific measurements ($\varepsilon$, ${H_{\text{eff}}}$) given a set of model parameters $\Omega_{\text{M}}$, $\Omega_{\Lambda}$, and $h$ (the curvature contribution is derived from $1=\Omega_{\text{M}}+\Omega_{\Lambda}+\Omega_{\text{K}}$). Fitting measurements with models amounts to sampling and maximizing $\cal{P}$ in parameter space. We use the emcee Markov-chain Monte Carlo (MCMC) package [117] to do the sampling, with multiple “walkers” mapping out the probability distribution $\cal{P}$ in the domain $\Omega_{\text{M}}=[0.045,1]$, $\Omega_{\Lambda}=[0,1]$, $h=[0.5,0.9]$. We use uninformative priors (all models in this domain are equally likely). We thus distribute the starting location of the walkers in parameter space with a low-discrepancy quasirandom number generator (scipy.qmc). We determine best-fits from the global maximum of the likelihood function as mapped by the MCMC walkers; we obtain uncertainties and correlations between parameters from the distribution of samples in parameter space.
@@
Figure 8: A corner of corner plots [118] showing probability distributions for parameter fits to void depth and outflow rates. The labels indicate the number of voids measured. The lines locate the true values of the parameters $h$ (orange), $\Omega_{\text{M}}$ (turquoise), and $\Omega_{\Lambda}$ (blue). The fits are most sensitive to $h$, then $\Omega_{\text{M}}$, and finally $\Omega_{\Lambda}$. In the upper corner plot, showing likelihood distributions for a single void, only the Hubble parameter is well constrained. Asserting prior knowledge of $h$ at its true value selects the probability distribution in a slice along the orange lines. With that prior, the (posterior) probability distribution for $\Omega_{\text{M}}$ becomes sharply peaked around its true value. In the lower left corner plot (two voids, $z=0$ and $z=1$), both $h$ and $\Omega_{\text{M}}$ are well constrained. In the lower right panel, with three voids ($z=0,0.5,1$), $\Omega_{\Lambda}$ is also well constrained.
@@ 4.2 The redshift range
The Hubble diagram built from Type Ia supernovae provides better constraints on cosmological parameters when applied over a large redshift range [119, 120]. Here, we show that voids similarly provide better constraints when we measure $fH$ as a function of $z$. Figure 9 illustrates this effect based on samples of 100 voids spaced uniformly in $z$. The broadest range of redshifts ($z=0$–2) yields the tightest constraints..
@@
There are two main challenges for measuring voids properties over a wide redshift range. At low redshift, the observable volume available to sample voids is small; in an absurd limit, there can be at most only one void of radius 100 $h^{-1}$pc within a survey of that depth. At high redshift the challenge is measurement of void properties from faint and distant galaxies. For this demonstration of the theoretical connection between void parameters and the underlying cosmology, we consider only the first challenge, identifying the number of voids we expect as a function of redshift.
Our approach for this assessment starts with an analytical estimate of the number of voids within a survey to a limiting redshift of $z=0.1$, a distance of roughly 300 $h^{-1}$pc in a $\Lambda$CDM universe. We consider voids with radii of 30–40 $h^{-1}$pc and depths $\varepsilon\geq 0.2$, corresponding to regions in the lowest quartile of the mass density distribution smoothed on the void size scale. Although a couple hundred of these voids could fit in the survey volume without overlapping, only about $n_{v}=50$ meet the depth criterion. Then, assuming that the comoving number density of voids is constant, we calculate the number of voids that lie in redshift bins of width 0.1 out to $z=2$. Figure 10 shows the results; the void counts rise steeply at low redshift ($z\lesssim 0.5$) and increase more slowly at higher $z$, reflecting the expansion of the background universe [121]. We also consider a void census that accounts for void evolution. At low redshift, 25% of the survey contain voids with $\varepsilon\geq 0.2$ as estimated from the statistics of the smoothed density field; at $z=2$, this fraction is only a few percent. Incorporating this factor when binning voids in figure 10, the binned void number counts exceed a thousand for $0.4\lesssim z\lesssim 2$, with a peak above 2,000 at $z\approx 1$.
@@ 4.3 Summary of cosmology from idealized voids
By measuring void depth $\varepsilon$ and outflow rate ${H_{\text{eff}}}$ in voids at multiple redshifts, we build samples of $fH$ that provide constraints on $h$, $\Omega_{\text{M}}$ and $\Omega_{\Lambda}$. This procedure mimics the approach to cosmological constraints based on Type Ia supernovae [116]. Preliminary tests with fits to measurements from solutions to full nonlinear shell model in section 3.3 indeed suggest that samples of void properties over a range of redshifts provide robust cosmological parameter estimates.
@@ 5 Voids in numerical realizations of the cosmic web
We create realizations of large chunks of the universe with an $n$-body particle algorithmm to evaluate the cosmological information encoded by voids. Our approach is based on the adhesion approximation [94]; similar $n$-body realization methods include the PATCHY code [122]. In our implementation, we use the Zel’dovich approximation in step-wise fashion, allowing particles to free stream with updates as large-scale structures emerge. Critically, the adhesion approximation adds small-scale viscous diffusion to mitigate the problem of trajectory crossings that arise when free streams of matter in the Zel’dovich flows pass right through one another. As in reference [94], this feature limits the spatial resolution of the realizations to about a megaparsec. Appendix A summarizes the implementation.
We adopt the adhesion approximation primarily because it is a fast, accurate way to track void dynamics. Because density fluctuations within voids are low amplitude, nonlinear dynamical effects are weak. Recent work [44, 40, 70] demonstrates that voids are well suited to the Zel’dovich approximation and the related adhesion algorithm. The realizations of large-scale structure we construct are geared toward toward modelling voids in the linear and translinear regimes; they are distinct from full $n$-body simulations that track small-scale nonlinear dynamics with Poisson solvers and ODE integrators.
@@ 5.1 Finding voids
A key step in assessing the connection between voids and their background cosmology is finding the voids. This process is complex mainly because the cosmic web is not Swiss cheese or even spongelike, with nicely spherical holes embedded in a high-density background. Instead, choices about density thresholds (voids are not vacuums), sizes and shapes are necessary and lead to a plethora of distinct void-finding algorithms, e.g., [18, 20, 109, 123, 124, 125, 52, 66, 67, 126, 127].
@@
We choose to implement our own void finding algorithm so that we could most easily assess the impact of changes in the algorithm parameters, balancing code speed, simplicity, and effectiveness at identifying comparatively shallow voids that fit with our analysis. Other options include the adoption of publicly available void finders [128, 129].
@@ 5.2 Cosmological parameter estimation
The theoretical curves in figure 14 are fingerprints of the background cosmology. To extract cosmological parameters by fitting the $n$-body samples shown in figure 14 with predictions from various cosmological models, we use the same MCMC-based parameter estimators as in section 4: We seek values in a parameter space of $\Omega_{\text{M}}$, $\Omega_{\Lambda}$ and $h$ using a $\chi$-square measure to assess the goodness of fit between theory (equation (3.16)) and samples of $(\varepsilon,{H_{\text{eff}}})$ from voids in the $n$-body realizations. Our priors are uniform inside the domain $1.050<\Omega_{\text{M}}<$ 1.05, and $-0.1<\Omega_{\Lambda}<1$, expanded slightly from the analysis of idealized voids to better explore the sensitivity of the estimators. The scatter in the samples used arises from deviations of the $n$-body realization voids from the ideal ones.
@@
Figure 15: Corner of corners plot [118] showing MCMC samples in parameter space for independent realizations of the $\Lambda$CDM model at redshifts from $z=0$ to $z=2$. Each realization lies in a cube of 500 $h^{-1}$pc to a side and has $n=512^{3}$ points. Voids of radii $30\pm 5$ $h^{-1}$pc are included. As in figure 9 for the ideal voids, the panels illustrate that a broader range of redshifts among void samples enhances parameter estimation.
@@ 5.3 The largest void in the universe
The distribution of void sizes can provide a powerful test of the background cosmology [21]. Here we focus on “the largest void”. The existence of extremely large voids may place strong constraints on the background model parameters [19]. The void identification criteria for this investigation are only slightly more stringent than finding local minima in tophat-smoothed density fields [24]. Voids in our catalog have profiles that are deep in their central region, and rise to a half-depth value at the nominal radius, behavior that requires specific phase alignment between density fluctuation modes. Supervoids — composed of multiple smaller voids — might be missed by our algorithm, depending on their substructure.
@@
Figure 16 shows an example of large voids within lightcone realization of 5000 $h^{-1}$pc to a side ($n=1024^{3}$). Because of grid effects, density modes are poorly sampled for voids with radii $\gtrsim$250 $h^{-1}$pc. All depths are scaled to the present epoch. For comparison, we also show the Eridanus Supervoid parameters [24]. The 1-$\sigma$ errors for the supervoid suggest that it is consistent with voids predicted in the $\Lambda$CDM realization. Since its initial discovery, the supervoid region has been reanalyzed [26, 27]. The supervoid is estimated to have a greater depth and smaller radial extent, but it is still consistent with standard theory.
Figure 16: Void depth versus size of the largest voids in a 5000 $h^{-1}$pc lightcone realization. This depth-size diagram applies to voids larger than 100 $h^{-1}$pc, scaled according to linear theory. The depth is effectively at the present epoch. The darker solid curve is $\sigma_{\text{R}}$ from the linear power spectrum; we eliminate void candidates with amplitudes below this value. The lighter-tone line is the amplitude of a one-of-a-kind void predicted in linear theory for the survey volume. The color of each sample point indicates whether the void is a “subvoid” (orange) or a void that is not embedded in a larger void (blue). The error bars show the observed Eridanus supervoid [24].
@@ 6 Voids in redshift space
So far, we have used the positions and peculiar velocities of galaxies as tracers to constrain the background cosmological parameters (section 5.2). However, observations do not yet measure accurate distances. Instead, extensive galaxy surveys provide the distribution of objects in redshift space. Here we consider the cosmological information content of voids as mapped in redshift space [130, 131, 46, 93, 132, 33, 133, 48, 97, 134].
@@
$\vec{s}$$\approx$ $\vec{S}+H\vec{r}+\hat{e}_{\text{los}}\cdot\vec{v}\hat{e}_{\text{los}}$ (6.1)
$\approx$ $\vec{S}+H\left(\vec{r}+\frac{f\varepsilon\Delta_{\text{los}}}{3}\hat{e}_{\text{los}}\right)$ (6.2)
@@
The mapping between real space and redshift space significantly impacts the measurement of void properties [106]. A small region in real space within an idealized void is stretched in redshift space along the line-of-sight by a factor of ${H_{\text{eff}}}/H$. The void thus appears deeper in redshift space:
$\varepsilon_{s}=\left(1+\frac{f}{3}\right)\varepsilon$ (6.3)
@@
Void shapes are also affected by redshift space distortions [106]. However, the net fractional elongation along the line-of-sight tends to be small because the outflow velocities are small compared with the Hubble flow across the void diameter. In the ideal case, we define the shape of a void in redshift space using the location of particles at the interface with the background universe. Neglecting any shell crossings, we approximate the void shape in redshift space as a prolate ellipse, with ellipticity
$e_{f}={H_{\text{eff}}}/H-1\approx\frac{f\varepsilon}{3},$ (6.4)
@@
Successful strategies to tease out the small-amplitude distortion of void geometry in redshift space focus on combining data from multiple voids. Aggregate measures like the void-galaxy correlation function quantify averaged redshift-space distortions that emerge away from the void centers, revealing peculiar motion [130]. For idealized voids, this information corresponds to the mean ellipticity at the void boundary. When calibrated with numerical simulations, the void-galaxy correlation function is impressively effective as a cosmological indicator [132, 48]. Outcomes include a measure of dark energy with the Alcock-Paczynski test and the velocity growth parameter $f$, as in references [132, 82, 97, 64].